OpenStudy (anonymous):

Part A: Solve A = 9 over 2(x + 28) for x. (4 points) Part B: Determine the value of x when A = 135. (2 points) Part C: Solve –np – 40 > 10 for n. Show your work. (4 points)

OpenStudy (anonymous):

i need help with this i dont understand how to do this

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

could you post a picture of the question?

OpenStudy (anonymous):

yea can you please post a picture I am kinda lost.

OpenStudy (anonymous):

that is the question

OpenStudy (anonymous):

9/2(x+28). solve for x

OpenStudy (anonymous):

thats what its asking

OpenStudy (anonymous):

i need help understanding how to solve for x

OpenStudy (anonymous):

use cross-multiplication

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

\[\frac{ 9 }{ 2(x+28) } = \frac{ 9 }{ 2X+56 }\]

OpenStudy (anonymous):

9 = a *2(x + 28)

OpenStudy (anonymous):

then leave x on one side and the rest on the other side and that's how you solve for x, i think ?

OpenStudy (anonymous):

Why did you put ''a'' in their?? @kx2bay

OpenStudy (anonymous):

im still not understanding?

OpenStudy (anonymous):

i didnt

OpenStudy (anonymous):

thats part of the question

hero (hero):

\[A = \frac{9}{2(x + 28)}\]

OpenStudy (anonymous):

That's the same thing I just put?

hero (hero):

To solve for x, you have to isolate x

hero (hero):

I would first re-write it as \[2(x + 28) = \frac{9}{A}\]

OpenStudy (anonymous):

so \[a=9 \over 2x+56?\]

hero (hero):

Then divide both sides by 2 Finally subtract 28 from both sides.

hero (hero):

@kryoWolf...no

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

hmm

hero (hero):

I've never seen a fraction written with the equal sign in it.

OpenStudy (anonymous):

i didnt know how not to put it like that in the equation like that

OpenStudy (anonymous):

@Hero how would \[9 \over A\] being divided by 2 look like?

OpenStudy (anonymous):

wait \[\frac{ x }{ 2 } +14=\frac{ 9 }{ A } \div 2?\]

hero (hero):

The 2 was already on the left side when you divided it out. You're misinterpreting what it means to divide that side by 2.

hero (hero):

The right side is correct.

OpenStudy (anonymous):

ok

hero (hero):

Here's what dividing both sides by 2 looks like: \(\dfrac{2(x + 28)}{2} = \dfrac{9}{2A}\)

OpenStudy (anonymous):

so \[x=\frac{ 9 }{ A }\div2-28?\]

OpenStudy (anonymous):

oh

hero (hero):

The 2's cancel on the left leaving just: \(x + 28 = \dfrac{9}{2A}\)

hero (hero):

Now subtract 28 from both sides to get: \(x = \dfrac{9}{2A} - 28\)

OpenStudy (anonymous):

ok

hero (hero):

So now, let A = 135, then calculate the value of \(x\)

OpenStudy (anonymous):

thanks for the help but i gotta go